. Fifty percent of Americans believed the country was in a recession, even though technically the economy had not shown two straight quarters of negative growth (BusinessWeek, July 30, 2001). For a sample of 20 Americans, make the following calculations. a. Compute the probability that exactly 12 people believed the country was in a recession. b. Compute the probability that no more than five people believed the country was in a recession. c. How many people would you expect to say the country was in a recession? d. Compute the variance and standard deviation of the number of people who believed the country was in a recession.


1
Expert's answer
2021-11-14T17:58:43-0500

Let X be the amount of american from sample of 20 people who believe country was in a recession.

Then X ~ Bin(20, 0.5)


(a) "P(X=12)={20 \choose 12}*(0.5)^{12}*0.5^8=0.1201"


(b) "P(X≤5)=P(X=0)+...+P(X=5)={20 \choose 0}*0.5^{20}+...+{20 \choose 5}*0.5^{20}=0.0000+0.0000+0.0002+0.0011+0.0046+0.0148=0.0207"


(c) We have to find the expected value of X

"E(X)=np=20*0.5=10" people we should expect


(d) "D(X)=np(1-p)=20*0.5*0.5=5" is the variance;

"\sigma(X)=\sqrt{D}=\sqrt{5}" is the standard deviation.